# Mistake Attribution: Fine-Grained Mistake Understanding in Egocentric Videos

### v1

@@ Abstract
We introduce Mistake Attribution (MATT), a task for fine-grained understanding of human mistakes in egocentric video. Unlike prior mistake understanding work, which lacks fine-grained output, MATT concretely attributes mistakes to the input instruction text or the attempt video. MATT determines what part of the instruction is violated (semantic role), when the deviation becomes irreversible (the Point-of-No-Return, PNR), and where the mistake appears in the PNR frame. We develop MisEngine, a data engine that automatically constructs attribution-rich mistake samples from existing datasets and inherits their annotations. Applied to large egocentric corpora, MisEngine yields EPIC-KITCHENS-M and Ego4D-M—two datasets that are up to two orders of magnitude larger than prior mistake datasets. We then present MisFormer, a unified attention-based model for mistake attribution across semantic (what), temporal (when), and spatial (where) dimensions, trained using MisEngine supervision. Experiments on our new datasets and prior benchmarks show that MisFormer outperforms strong video-language, temporal localization, hand-object interaction, and mistake-detection baselines.
@@ 1 Introduction
Physically-grounded instructional AI assistants that provide guidance and feedback given egocentric video has drawn significant attention in the vision community [2, 46, 42]. Leveraging clear hand and object interaction [41, 20] along with rich task models [30], these assistants show significant promise in the way humans acquire skills [6, 42] across various physical activities, spanning everyday activities like cooking, cleaning, and maintenance [53, 12, 2, 40, 32, 46] to professional tasks [13, 11, 55, 35, 3].
@@ 2 Related Work
Modeling Mistakes in Egocentric Videos. Although there is a large body of early work in video anomaly detection [33, 50, 52, 54, 14], most of this work does not emphasize anomalies as mistakes in the context of a procedure. Most past mistake-specific work focuses on detection or small-set categorization. Soran et al. [43] identify omitted actions using a hidden Markov models. HoloAssist [46] uses the TimeSFormer ViT model [1, 5] to detect mistakes. Assembly101 [40, 4] uses a graph-based method to detect two categories of mistakes specifically in the assembly task: the ordering and placement of installing a component. Certain recent works also adopt the anomaly detection mindset for mistakes, such as EgoPED [23], which compares an observed action feature to its prototypical one, PREGO [9], which compares the inferred action class to the expected one in a streaming manner, AMNAR [16], which combines these two ideas, and Missteps [31], which compares observed gaze behavior against an expected one. None of these works models mistakes at a level of richness beyond binary detection.
The most relevant recent works are twofold. Lee et al. [22] categorize detected mistakes into a set of predefined categories (e.g., modification, slip errors, etc.). MistSense [36] generates natural language explanations for an action video that is detected as mistaken. Neither of these works captures a sufficiently rich task-based analysis of the mistake. The categories in Lee et al. [22] do not describe the specific attributes of the mistake. MistSense’s [36] explanation capture generic action failures (e.g., knocking over the bottle) rather than how the action deviates from the task instructions.
@@ 4.1 Setup
We benchmark attribution subtasks on Ego4D-M and EPIC-KITCHENS-M. We split each dataset into training/validation/test sets in an 8:1:1 ratio. This yields 206K/25K/25K for Ego4D-M and 176K/22K/22K samples for EPIC-KITCHENS-M. We additionally use EgoPER [23], a popular existing mistake dataset, for mistake detection evaluation.
@@ 4.2 Results
Hand-detection methods fail to address the spatial attribution task because they do not localize the manipulated object, which is important for precise mistake grounding and downstream recovery [26]. MisFormer substantially outperforms hand-detection baselines (MediaPipe-U) for spatial attribution on the PNR frame, improving mIoU by +9.33 and reducing Center Distance (CD) and Box Size Error (BSE) by 3.11 and 4.71, respectively. Compared to a SOTA hand-object interaction detector (SSDA), MisFormer is less accurate on localization (mIoU 59.21 vs. 68.54; CD 10.36 vs. 4.52; BSE 16.27 vs. 8.56). Nevertheless, MisFormer offers a unified framework which performs semantic and temporal attribution in conjunction with localization. MisFormer is also efficient. Considering only the spatial attribution module, it runs at 68.9 FPS versus 28.44 FPS (SSDA) on the same NVIDIA H100 GPU.

### v2

@@ Abstract
We introduce Mistake Attribution (MATT), a new task for fine-grained understanding of human mistakes in egocentric videos. While prior work detects whether a mistake occurs, MATT attributes the mistake to what part of the instruction is violated (semantic role), when in the video the deviation becomes irreversible (the Point-of-No-Return, PNR), and where the mistake appears in the PNR frame. We develop MisEngine, a data engine that automatically constructs mistake samples from existing datasets with attribution-rich annotations. Applied to large egocentric corpora, MisEngine yields EPIC-KITCHENS-M and Ego4D-M—two datasets up to two orders of magnitude larger than prior mistake datasets. We then present MisFormer, a unified attention-based model for mistake attribution across semantic, temporal, and spatial dimensions, trained with MisEngine supervision. A human study demonstrates the ecological validity of our MisEngine-constructed mistake samples, confirming that EPIC-KITCHENS-M and Ego4D-M can serve as reliable benchmarks for mistake understanding. Experiments on both our datasets and prior benchmarks show that MisFormer, as a single unified model, outperforms task-specific SOTA methods by at least 6.66%, 21.81%, 18.7%, and 3.00% in video-language understanding, temporal localization, hand-object interaction, and mistake detection, respectively.
@@ 1 Introduction
Physically-grounded instructional AI assistants that provide guidance and feedback given egocentric video has drawn significant attention in the vision community [5, 53, 47, 2]. Leveraging clear hand and object interaction [46, 23] along with rich task models [35], these assistants show significant promise in the way humans acquire skills [9, 47] across various physical activities, spanning everyday activities like cooking, cleaning, and maintenance [60, 15, 5, 45, 37, 53] to professional tasks [16, 14, 62, 40, 6].
@@ 2 Related Work
Modeling Mistakes in Egocentric Videos. Recent efforts toward Instructional AI have demonstrated early progress [2, 30, 49, 31, 1, 52, 4]. In terms of Misatake Understanding, although there is existing work in video anomaly detection [38, 57, 59, 61, 17], most of this work does not emphasize anomalies as mistakes in the context of a procedure. Most past mistake-specific work focuses on detection or small-set categorization. Soran et al. [48] identify omitted actions using a hidden Markov models. HoloAssist [53] uses the TimeSFormer ViT model [3, 8] to detect mistakes. Assembly101 [45, 7] uses a graph-based method to detect two categories of mistakes specifically in the assembly task: the ordering and placement of installing a component. Certain recent works also adopt the anomaly detection mindset for mistakes, such as EgoPED [26], which compares an observed action feature to its prototypical one, PREGO [12], which compares the inferred action class to the expected one in a streaming manner, AMNAR [19], which combines these two ideas, and Missteps [36], which compares observed gaze behavior against an expected one. None of these works models mistakes at a level of richness beyond binary detection.
The most relevant recent works are twofold. Lee et al. [25] categorize detected mistakes into a set of predefined categories (e.g., modification, slip errors, etc.). MistSense [41, 49] generates natural language explanations for an action video that is detected as mistaken. Neither of these works captures a sufficiently rich task-based analysis of the mistake. The categories in Lee et al. [25] do not describe the specific attributes of the mistake. MistSense’s [41] explanation capture generic action failures (e.g., knocking over the bottle) rather than how the action deviates from the task instructions.
@@ 4.1 Setup
We benchmark attribution subtasks on Ego4D-M and EPIC-KITCHENS-M. We split each dataset into training/validation/test sets in an 8:1:1 ratio. This yields 206K/25K/25K for Ego4D-M and 176K/22K/22K samples for EPIC-KITCHENS-M. We additionally use EgoPER [26], an existing mistake dataset, for mistake detection evaluation.
@@ 4.2 Results
Hand-detection methods fail to address the spatial attribution task because they do not localize the manipulated object, which is important for precise mistake grounding and downstream recovery [29]. MisFormer substantially outperforms hand-detection baselines (MediaPipe-U) for spatial attribution on the PNR frame, improving mIoU by +9.33 and reducing Center Distance (CD) and Box Size Error (BSE) by 3.11 and 4.71, respectively. Compared to a SOTA hand-object interaction detector (SSDA), MisFormer is less accurate on localization (mIoU 59.21 vs. 68.54; CD 10.36 vs. 4.52; BSE 16.27 vs. 8.56). Nevertheless, MisFormer offers a unified framework which performs semantic and temporal attribution in conjunction with localization. MisFormer is also efficient. Spatial attribution module runs at 68.9 FPS versus 28.44 FPS (SSDA) on the same NVIDIA H100 GPU.
@@ Conclusion
Acknowledgement
This research was funded, in part, by the U.S. Government under ARPA-H contract 1AY2AX000062. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the U.S. Government.
Supplementary Material
A MisEngine Details
Role-matching function $\mathcal{J}$. $\mathcal{J}(g^{r}_{i},g^{r}_{j})$ determines whether role $r$ matches between two action descriptions, and thus whether a cross-matched video contains a mistake on that role. The matching rule is granularity-dependent: e.g., “cut” vs. “dice” are distinct in a professional kitchen but interchangeable in casual cooking. We implement two variants(Sec. 3.1): (1) taxonomy-based $\mathcal{J}$ for Ego4D, grouping synonymous verbs and nouns from the original dataset; and (2) literal string matching for EPIC-KITCHENS. SRL on 1K Ego4D descriptions yields 97.97% predicate and 94.13% object F1, reflecting the short imperative phrasing of the source annotations.
B Dataset Ecological Validation
[…]
EK-M (ours) Ego4D-M (ours) EgoPER [23] CC4D [37]
[…]
(a) % of samples human-rated as realistic (Real-% $\uparrow$) on mistake videos.
Tmp. (%) Spa. (%)
96.29 $\pm$ 3.99 95.81 $\pm$ 2.16
(b) Real-% on inherited annotations.
$\text{Mis}_{V}$ $\text{Mis}_{Obj}$ $\text{Mis}_{Both}$
60.73 57.99 58.91
(c) Spatial-Attribution mIoU by mistake type.
Tab. (a): All four datasets receive comparably high realism ratings (within std). This is expected since cross-matched samples are drawn from the same domain, grounded in original metadata. Participants occasionally flag samples in prior datasets as “staged,” consistent with the manual collection challenges discussed in Sec. 1.
Tab. (b): Participants judge whether the inherited PNR frame marks the consolidation of a mistake (Tmp.) and whether the bounding box captures mistake-relevant detail (Spa.). Both achieve >>95% Real-%, validating the quality of inherited annotations. The temporal attribution frame occurs at 79.4879.48% of video duration on average. Samples are constructed in equal numbers across mistake types for balanced training; this count is configurable (Sec. 3.1).
C Additional Experiments
Zero-shot transfer to EgoPER. MisFormer generalizes zero-shot to EgoPER, achieving $34.46$ F1 / $35.52$ Acc and surpassing both baselines. Tab. 5 results show that manually collected datasets provide insufficient training signal for MisFormer (Tab. 1), underscoring the necessity of MisEngine.
Spatial attribution mIoU by mistake type. Tab. (c) reports spatial attribution mIoU broken down by mistake type. $\text{Mis}_{\text{Both}}$ (which uses a union bounding box over hand and object) performs comparably to single-role types. Union boxes remain compact because, at the PNR frame, hand and object regions substantially overlap during active interaction.
Ablation: semantic head and projection block. Tab. 6 in the main paper ablates the projection block $\mathcal{P}$ and task-specific heads. Removing $\mathcal{P}$ degrades all tasks (row c). Removing temporal or spatial supervision each hurts the complementary task (rows d, e), confirming their mutual benefit. We additionally ablate the semantic head: without it, semantic attribution and mistake detection are disabled, while temporal and spatial attribution improve to $0.358$ s MAE and 65.3365.33% mIoU. This suggests the current $\mathcal{P}$ faces tension between representing semantic-level and grounded-level features simultaneously, which we identify as a direction for future work. The unified structure nonetheless avoids training separate models, runs $2.42\times$ faster (Sec. 4.2), and outperforms 6 of 7 baselines (Tabs. 2, 4 and 3).
D Qualitative Results and Failure Analysis
Qualitative examples. Fig. 5 shows representative outputs. Row 1 correctly produces semantic attribution, localizes the PNR frame, and generates a reasonable spatial box (IoU=0.56). Row 2 illustrates a failure case: semantic attribution is correct, but temporal attribution selects an earlier frame before the bag is set down. Judging the spatial completion state of an object from 2D video alone is inherently difficult, suggesting that depth or 3D cues are a promising direction for future work. Notably, the spatial bounding box still correctly highlights the hand performing the unintended action, which is consistent with the correct semantic prediction.
Figure 5: Top: correct predictions (IoU=0.56). Bottom: failure case with correct semantic but incorrect temporal attribution.
E Implementation Details
[…]